• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Sirca

    Log entry time 21:55:54 on March23,2005

    Entry number 142265

    keyword=Beam being retuned in injector and linac

    The MCC can not tune the beam well. It looks bad (large halo) not
    only at the Compton location, but also in the injector and linac.
    And they are having simultaneous, related problems with Hall B.
    They will thus try to retune the injector and linac sections,
    then try to go through Compton; if the resulting beam at our target
    still looks reasonable for SRC in terms of harp scans, we will take
    it and abandon the chicane. This is agreed upon with Sirish.